### v1

# Adaptive Agents in Spatial Double-Auction Markets: Modeling the Emergence of Industrial Symbiosis

Abstract.
Industrial symbiosis fosters circularity by enabling firms to repurpose residual resources, yet its emergence is constrained by socio-spatial frictions that shape costs, matching opportunities, and market efficiency. Existing models often overlook the interaction between spatial structure, market design, and adaptive firm behavior, limiting our understanding of where and how symbiosis arises. We develop an agent-based model where heterogeneous firms trade byproducts through a spatially embedded double-auction market, with prices and quantities emerging endogenously from local interactions. Leveraging reinforcement learning, firms adapt their bidding strategies to maximize profit while accounting for transport costs, disposal penalties, and resource scarcity. Simulation experiments reveal the economic and spatial conditions under which decentralized exchanges converge toward stable and efficient outcomes. Counterfactual regret analysis shows that sellers’ strategies approach a near Nash equilibrium, while sensitivity analysis highlights how spatial structures and market parameters jointly govern circularity. Our model provides a basis for exploring policy interventions that seek to align firm incentives with sustainability goals, and more broadly demonstrates how decentralized coordination can emerge from adaptive agents in spatially constrained markets.
Key words and phrases:
Agent Based Simulation, Industrial Symbiosis, Circular Economy, Market Analysis
1. Introduction
Industrial Symbiosis (IS) refers to the exchange of byproducts between firms, turning one company’s waste into valuable inputs for another one. By reusing materials such as heat, rubble, or chemical residues, IS reduces environmental impact while generating economic benefits, embodying the circular economy principle of closing resource loops. For instance, the RETDA eco-industrial park in China reported a 33% increase in input productivity, a 3650% rise in water use efficiency, and a 30.91% reduction in emissions chen2022.
Despite its potential, IS rarely emerges spontaneously. Exchanges depend on local supply-demand compatibilities, economic incentives, and logistical constraints henriques2021industrial. Poor coordination often results in wasted resources and missed opportunities, highlighting the gap between individual strategies and collective circularity goals. Identifying the conditions under which decentralized firms can coordinate effectively is therefore a key research challenge.
Analytical solutions are often intractable due to the combinatorial complexity of interactions and the emergent nature of prices and traded quantities. Simulation-based approaches, by contrast, are particularly suited to study such nonlinear, decentralized, and heterogeneous systems he2019heterogeneous. By systematically varying system configurations, simulations can indeed reveal market behaviors and highlight factors that support efficient and resilient local circularity. Agent-based models, in particular, are able to capture heterogeneity among actors and test how different market designs, decision rules, and learning mechanisms affect the emergence of exchanges under realistic economic constraints epstein1996growing. Their ability to explicitly represent geosituated agents makes them especially suitable for studying spatially dependent interactions, such as proximity based collaborations, transportation costs, and localized resource flows. Unlike theoretical game-theoretic models, which often rely on a small number of representative players to ensure tractability, agent-based models can accommodate many heterogeneous actors and interactions.
However, their effectiveness is often constrained by the simplifications made in representing agent behaviors. In many existing symbiosis models Yazan2020; Ghali2017; Raimbault2020; Gan2015; FernandezMena2020; Chahla2019, firms follow fixed behavioral rules or static pricing strategies, limiting their ability to capture the adaptive and uncertain nature of real-world exchanges. In practice, firms continuously adjust decisions in response to shifting demand, fluctuating costs, and evolving environmental regulations. This highlights the need to enhance decision-making mechanisms within agent-based models, enabling agents to learn, adapt, and explore strategies in a more dynamic way.
In this paper, we present a spatially explicit, decentralized multi-agent model in which buyers and sellers of byproducts interact through multilateral spatial double auctions using co-evolving strategies. Sellers adaptively update their pricing strategies via reinforcement learning in a partially observable environment, while buyers select offers that maximize utility subject to spatial and economic constraints. A key feature of the model is that both transaction prices and traded quantities emerge endogenously from these interactions, allowing us to investigate how local decisions give rise to systemic patterns. Our contributions are threefold. First, we develop a reinforcement learning-based market model for decentralized byproduct exchanges, based on a novel auction mechanism adapted to spatial and economic constraints of territorial circularity. Second, we validate the learned strategies via counterfactual regret analysis, showing convergence toward near-Nash equilibria. Finally, a simulation-based analysis of conditions under which adaptive pricing and agent interactions lead to stable, efficient, and circular outcomes, providing insights for policymakers.
The paper first reviews related work (Section 2), then presents the decentralized market model (Section 3) and simulation setup (Section 4). Results on price formation and local circularity are analyzed in Section 5, followed by discussion (Section 6) and conclusions (Section 7).
2. Related Work
2.1. Industrial Symbiosis and Circular Economy Modeling
Industrial symbiosis in Eco-Industrial Parks (EIPs) has traditionally been studied using system dynamics to capture network level behavior and feedback loops zhao2008simulation; qu2010system; sterman2000business; lorenz2006towards; cao2009applying. Complementary approaches have treated EIPs as complex adaptive systems, highlighting dynamic business interactions and emergent symbiotic relationships chertow2012organizing; yu2013process; felicio2016; lambert2002eco; romero2013framework. Other studies focused on material and energy flows, byproduct exchanges, and social network analysis Romero2014; felicio2016; yazan2016design; chertow2000; domenech2009social. Recent reviews Demartini2022; rizzati2024systematic emphasize the potential of agent-based modeling to capture local, adaptive, and self-organizing behaviors. Emerging agent-based studies have integrated input-output models Yazan2020, explored spatially networked exchanges Ghali2017; Raimbault2020; mastioCIEC, and examined sector specific synergies Gan2015; FernandezMena2020; Chahla2019. Yet, despite these advances, the explicit representation of economic incentives and market coordination mechanisms remains underexplored. Also, these works does not explicitly model markets where prices, trade volumes, and resource allocations emerge endogenously from decentralized interactions.
2.2. Auctions and Market Mechanisms in Spatial Settings
Auction and negotiation mechanisms in Multi-Agent Systems (MAS) have become canonical tools for decentralized coordination. From early negotiation platforms Chavez1997 to double auctions—market mechanisms in which multiple buyers and sellers simultaneously submit bids and offers—used to study bidding strategies and price discovery Tesauro2001; Satterthwaite2022, research has demonstrated how decentralized negotiations can generate robust outcomes. Extensions have examined repeated competition Ledvina2010; Bichler2024, peer-to-peer trading Qiu2021, social influence wainwright2024social, and multilateral bargaining across multiple resources Chen2015; An2011. Spatial aspects are frequently acknowledged, for instance through costly trade links babaioff2005, heterogeneous bilateral transport costs chu2008, or land markets where localized interactions induce differentiated prices filatova2007. Yet these contributions remain largely theoretical, typically involving few participants and limited attention to adaptive mechanisms that capture the evolving dynamics of real markets. In particular, the integration of spatial frictions with learning-based bidding strategies has received limited attention, despite its relevance for understanding decentralized coordination in resource-constrained environments.
2.3. Learning Agents for Adaptive Market Participation
Traditional learning agents often rely on fixed rules Cliff2024, or supervised training lauretto2013evaluation, limiting their ability to adapt to dynamic market conditions and strategic interactions. Reinforcement Learning (RL) overcomes these limitations by allowing agents to learn optimal policies through trial-and-error, balancing exploration and exploitation in complex, evolving environments. In the context of double auctions, RL agents are particularly well suited due to the complex interactions between buyers and sellers and the continuous feedback provided by order book dynamics. Agents can be designed to adopt specific roles, such as deciding when to buy, sell, or set prices, enabling cooperative and strategic trading behaviors bjerkoey2019; maeda2020. Multi-Agent Reinforcement Learning (MARL) extends this approach to settings with multiple interacting agents, allowing the simulation of bottom-up market dynamics shavandi2022; karpe2020; lussange2020; ganesh2019. MARL models have been shown to reproduce emergent behaviors in order books and centralized market microstructures by capturing interactions among heterogeneous agents karpe2020. This enables detailed analysis of price formation, liquidity, and strategic adaptation. However, existing approaches largely ignore spatial constraints, as these are typically absent in the stock market. Yet, such constraints can be critical in markets where transportation or delivery costs influence trading decisions rodrigue2020geography. RL with spatial considerations have been primarily explored in logistics, where carriers and shippers act as bidders in dynamic auctions mediated by centralized brokers Wang2018; Qiao2019; zha2017surge; atasoy2020. Even when shippers themselves are modeled as learning agents VanHeeswijk2019, exchanges remain intermediary mediated. In contrast, many real-world IS transactions are inherently local and broker-free. To the best of our knowledge, RL has not yet been applied to decentralized, spatially constrained markets where transport costs and policy incentives influence agents’ strategies.
2.4. Positioning of This Work
Taken together, three main research gaps emerge:
  1. (1)
    Industrial symbiosis models rarely provide in depth mechanism for endogenous market-based price formation.
  2. (2)
    Auction and negotiation models seldom incorporate adaptive learning of agents.
  3. (3)
    Reinforcement learning in markets has not been applied to decentralized settings where spatial topology influence agents behavior.
Our contribution is to bridge these strands by introducing a decentralized auction model for industrial symbiosis where (i) sellers adapt prices through reinforcement learning, (ii) spatial constraints are explicitly internalized, and (iii) both prices and traded quantities emerge from multilateral interactions. This framework connects industrial symbiosis modeling with mechanism design and multi-agent RL, allowing modelers to explore how local adaptive dynamics shape circularity outcomes. Our objective is to reproduce the systemic complexity of inter-firm interactions within industrial symbiosis as a subset of the broader circular economy, providing a foundational basis for modeling circular economic dynamics.
3. Model
3.1. Environment and Agents
To concentrate on the essential dynamics of the model, we restrict our analysis to the exchange of a single byproduct. This simplification enables us to isolate and assess the effects of multilateral interactions and sellers strategies, without additional complexity introduced by multiple interdependent byproducts. We model a decentralized market for one byproduct, composed of buyers $\mathcal{B}=\{B_{1},\dots,B_{N_{B}}\}$ and sellers $\mathcal{S}=\{S_{1},\dots,S_{N_{S}}\}$. Each agent has a spatial location, and the distance between a buyer $B_{i}$ and a seller $S_{j}$ is denoted $d_{ij}$. Buyers have heterogeneous initial demands $q_{i}$ drawn from a scaled uniform distribution and evaluate incoming offers based on their price sensitivity $\beta_{i}$. Specifically, buyer $B_{i}$ accepts an offer $p_{ij}$ from seller $S_{j}$ only if $p_{ij}\leq\beta_{i}\,p_{m}$, where $p_{m}$ is the reference market price. This mechanism ensures that buyers only engage in transactions that meet their cost expectations relative to the external market.
Sellers are endowed with quantities $q_{j}$ and propose personalized prices to buyers, defined as
$p_{ij}=\phi_{j}\ p_{m}+c_{ij},$ (1)
The term $c_{ij}$ is originally defined as $c_{ij}=d_{ij}\ c_{t}$, where $c_{t}$ represents the transportation cost per kilometer (expressed in percentage of the market price). This term can be further extended to capture the total logistics cost required for the byproduct provided by the seller to become usable by the buyer, including handling, treatment, or adaptation costs. Finally, $\phi_{j}$ is an adaptive pricing parameter, that evolves during the learning phase.
In addition to market and transportation costs, sellers face a penalty $c_{d}$ for unsold quantities (also expressed in percentage of the market price), representing the cost of sending residual flows to landfill and reflecting both economic and environmental impacts. If a seller $j$ ends a timestep with $q_{j}^{\text{unsold}}$ units remaining, its reward is reduced by $q_{j}^{\text{unsold}}\ c_{d}$, imposing a direct financial penalty that encourages sellers to adjust pricing strategies to minimize waste while still aiming to maximize profit.
To characterize market conditions, we define the scarcity level of the byproduct as
$s=\frac{\sum_{i\in\mathcal{B}}q_{i}}{\sum_{j\in\mathcal{S}}q_{j}}.$ (2)
This measure allows us to systematically investigate the effects of supply-demand imbalances on agents’ behavior in the decentralized market.
Each seller $S_{j}$ aims to maximize its cumulative expected profit at each simulation step. Let $\mathbf{q}_{j}^{t}=(q_{ij}^{t})_{i\in\mathcal{B}}$ denote the vector of quantities bought by $B_{i}$ from $S_{j}$ at step $t$, and $\mathbf{p}_{j}^{t}=(p_{ij}^{t})_{i\in\mathcal{B}}$ the corresponding price vector. Defining $\mathbf{c}_{j}=(c_{ij})_{i\in\mathcal{B}}$ as the transport cost vector, the instantaneous profit can be written compactly as
$r_{j}^{t}=\langle\mathbf{q}_{j}^{t},\,\mathbf{p}_{j}^{t}-\mathbf{c}_{j}\rangle\;-\;q_{j}^{\text{unsold},t}\ c_{d},$ (3)
where $\langle\cdot,\cdot\rangle$ denotes the dot product. The first term corresponds to the net revenue from executed contracts, while the second term penalizes unsold inventory through the landfill cost $c_{d}$. This spatially defined market closely resembles a spatial Bertrand oligopoly hotelling_stability_1929, where multiple sellers compete by setting prices and buyers prefer lower-cost options. Unlike the classical Bertrand model, where identical products drive prices toward marginal cost, the inclusion of transport costs differentiates sellers spatially, allowing local price variation and creating room for adaptive strategies to emerge.
3.2. Decentralized Multilateral Auction
We model the interactions between buyers and sellers using a decentralized multilateral auction described in Algorithm 1. Each seller submits a set of bids to each buyers, where each bid specifies a price and quantity for a given product $p$, computed on the basis of the logistic costs required to deliver the product to the buyer’s location and the external market price. Buyers evaluate the incoming bids and select the one that maximizes their utility, subject to their price sensitivity $\beta_{i}$ and demand $q_{i}$. Selected bids are then sent back as proposals to sellers. Each seller chooses a subset of the received bids to accept in order to maximize expected profit, taking into account transport costs $c_{ij}$ and remaining inventory $q_{j}$. Contracts are executed, updating both buyers’ needs and sellers’ available quantities. This process iterates until no further mutually acceptable bids can be matched. Sellers adapt their pricing strategy $\phi_{s}$ over time using a strategy based on historical rewards that will be described in 3.3.
Algorithm 1 Decentralized Multilateral Auction
1:Buyers $\mathcal{B}$, Sellers $\mathcal{S}$, product $p$
2:while contracts has been made at the previous timestep do
3:  Seller bids: each $S_{j}\in\mathcal{S}$ submits a bid $(q_{j},p_{ij})$ to each $B_{i}\in\mathcal{B}$, where $q_{j}$ is the maximum quantity available for $S_{j}$, and $p_{ij}=\phi_{j}\ p_{m}+d_{ij}\ c_{t}$
4:  Buyer choice: each $B_{i}$ evaluates the received bids, selects the most attractive one within price tolerance $\beta_{i}p_{m}$, and sends acceptance proposal to the corresponding seller
5:  Seller allocation: each $S_{j}$ sorts accepted bids by expected profit per sold unit, confirms contracts until inventory is exhausted, updates sold quantity and profit
6:  Remove satisfied buyers and empty sellers from $\mathcal{B}$ and $\mathcal{S}$
7:end while
8:Return executed contracts
3.3. Seller Reinforcement Learning and Action Selection
It has been shown by Bichler2024 that Bertrand competition problems can be effectively learned using bandit algorithms. Accordingly, each seller’s sequential pricing decision is modeled as a contextual multi-armed bandit. The action space is discretized into $K$ intervals, $\forall\ 0<k\leq K$:
$\Phi=\{\phi^{(1)},\dots,\phi^{(K)}\},\quad\phi^{(k)}=\phi_{\min}+\frac{k-1}{K-1}(1-\phi_{\min}),$ (4)
where $\phi_{\min}=-\frac{c_{d}}{p_{m}}$. At every timestep $t$, seller $S_{j}$ selects a discrete pricing action $\phi_{j}^{t}\in\Phi$ based on the current market context and receives a reward $r_{j}^{t}$. Each interval is associated with a weight $w_{k}$ representing the expected reward. When an action is chosen and the associated reward is obtained, the weights are updated via an exponential moving average (EMA):
$w_{k}^{(t+1)}=\begin{cases}\alpha w_{k}^{t}+(1-\alpha)r_{t},&\text{if }\phi_{j}=\phi^{(k)},\\ w_{k}^{t},&\text{otherwise}.\end{cases}$ (5)
Here, $\alpha\in[0,1]$ is the smoothing factor, which balances the influence of past and recent observations. By prioritizing recent data, EMA allows rapid adaptation to changing conditions while remaining computationally efficient, requiring only constant-time updates per action. Although it does not offer formal convergence guarantees, EMA provides a practical compromise between adaptivity and scalability, and allows to maintain high learning efficiency morales2024exponential.
Sellers select their next action using Boltzmann (softmax) sampling sutton1998reinforcement with a declining temperature $\tau_{t}$:
$P_{k}=\frac{\exp(w_{k}/\tau_{t})}{\sum_{l=1}^{K}\exp(w_{l}/\tau_{t})},\quad\tau_{t}=\max(\tau_{\min},\tau_{0}\cdot\text{decay}^{t}),$ (6)
where $\tau_{0}$ sets the initial exploration level, decay controls the annealing rate, and $\tau_{\min}$ prevents premature convergence. This ensures exploration dominates early in learning, while the policy gradually becomes greedier over time. The subsequent action $\phi_{j}^{(t+1)}$ is drawn at random from the set of actions according to probabilities $P_{k}$. This sampling strategy balances exploration and exploitation in a multi-agent, stochastic environment. In our simulations, Boltzmann exploration consistently converged faster and with less oscillation than $\epsilon$-greedy or UCB, which often struggled with noise from multi-agent interactions o2021variational. By combining a discretized action space, adaptive reward estimates, and Boltzmann sampling, the model captures the decentralized emergence of price formation and allocation dynamics in a spatially distributed market for byproducts. While Boltzmann does not always guarantee optimal regret bounds in multi-agent stochastic settings, its smooth adaptation makes it practical and effective for this scenario.
3.4. Symbiosis Index
We finally introduce a symbiosis index (SI) to quantify the efficiency of local byproduct exchanges. It is defined as:
$\text{SI}=\frac{q_{\text{bought}}}{\min({q_{\text{toSell}},q_{\text{needed}}})},$
where $q_{\text{bought}}$ denotes the quantity actually purchased, $q_{\text{toSell}}$ the quantity offered by the seller, and $q_{\text{needed}}$ the quantity required by the buyer. The denominator normalizes the indicator so that it cannot exceed 1, even when the sold quantity is greater than the buyer’s need. This ensures that the symbiosis index reflects the proportion of demand satisfied rather than rewarding oversupply. In contrast to existing technical indicators hardy2002industrial; tiejun2010two; zhao2008simulation; gao2013study, which often incorporate numerous parameters (e.g., material compatibility, lifecycle impacts, or multi-criteria environmental scores), our simplified formulation deliberately focuses only on exchanged quantities. This avoids unnecessary complexity while remaining sufficient for capturing efficiency in the present modeling framework.
4. Experiments
4.1. Simulation Framework
The model is implemented in a simulation framework that supports decentralized markets with heterogeneous agents. Buyers and sellers are spatially distributed across a territory, which can be represented either with an abstract 2D environment or with a real geographic area with road network constraints. Distances, travel times, and transport costs are computed based on the underlying spatial representation, allowing realistic modeling of logistical constraints. The businesses interact through the auction mechanism described earlier, with offers, acceptances, and executed contracts dynamically updating inventories and demands.
The framework tracks a variety of systemic metrics over time, including:
  • Local prices and their evolution;
  • Proportion of demand satisfied locally (symbiosis index);
  • Network connectivity and trading patterns;
  • Distance travelled by the sold byproducts.
These indicators form the backbone of our analysis, as they connect local agent decisions to measurable systemic dynamics. This flexibility enables us to investigate a wide range of experimental settings, from dense and highly interconnected territories to sparse environments where exchanges are scarce. Visualization tools provide spatial and temporal representations of market dynamics, price formation, and circularity indicators. The code implementation is available on GitLab [^1] https://github.com/AAMAS746/aamas2026-746.
4.2. Experimental Conditions.
For similar reasons behind our decision to focus on a single product, we refrain from using a real-world territory and instead construct virtual environments. This approach provides precise control over key parameters such as firm density, territorial extent, and spatial clustering, thereby enabling a systematic exploration of their influence on circularity indicators and overall market performance. One of the key factors shaping the potential for cooperation is the spatial dispersion of firms. We therefore define the cluster spread parameter ($cs$) as the standard deviation of companies positions around cluster centers, normalized by the environment width: small values lead to compact, well separated clusters, while setting it to $1$ produces a configuration comparable to a fully uniform distribution across the environment, effectively removing any clustering effect. Figure 1 shows the difference between a low (1(a)) and high (1(b)) spread value. In our experiments, businesses were allocated into four spatial clusters. By default, the cluster spread parameter was set to $1$, indicating a broad dispersion and minimal clustering.
(a) $cs=0.05$
(b) $cs=0.5$
Figure 1. Comparison of spatial agent distributions for different cs values.
The results we obtain came from simulations conducted with a population of 40 firms. The market price was fixed at $100$, and the transportation cost was set to $0.1$ per kilometer. Sellers adjusted their strategies through Boltzmann exploration, with a temperature decay parameter set to $0.996$ and the action space discretized into $30$ bins. The relatively small scale of the simulated market was chosen to balance computational tractability with the ability to capture complex multi-agent interactions. This size of population allows us to conduct thousands of simulation runs for systematic sensitivity analyses and counterfactual experiments, which would be computationally prohibitive at larger scales. Furthermore, each simulation runs for a large number of timesteps (1000 steps) to ensure that seller strategies have sufficient time to adapt and converge, and that emergent patterns of price formation and allocation stabilize. By prioritizing repeated experimentation and long-horizon dynamics over absolute market size, we focus on understanding the fundamental mechanisms driving spatial effects, adaptive bidding behavior, and the emergence of circularity. Our implementation is optimized for performance, allowing on a standard laptop processor (Intel(R) Core(TM) Ultra 9 185H processor with 21 cores) to achieve approximately 100 simulations of 1000 timesteps per minute.
5. Results
In this section, we examine the core dynamics of the decentralized, spatially explicit market, focusing on conditions for price convergence and the efficiency of adaptive strategies using counterfactual regret analysis. We then explore how spatial organization, resource scarcity, disposal costs, and density interact to shape market outcomes and local circularity. Finally, a global sensitivity analysis quantifies the influence of individual parameters and their interactions on emergent symbiosis patterns.
5.1. Convergence Toward Equilibrium
We first explore empirically the evolution of transaction prices as sellers adapt their strategies under different scarcity and disposal cost conditions, with a density fixed to 0.001. Figures 2(a) and 2(b) show the mean price trajectories over time compared to the external market price and the disposal cost. As anticipated given the decaying temperature, the curves gradually stabilize, approaching an equilibrium after sufficient simulation steps. In the high scarcity and low disposal cost setting (Figure 2(a)), prices converge toward an equilibrium close to the external market price. By contrast, in the low scarcity and medium disposal cost scenario (Figure 2(b)), prices decline steadily and stabilize near the disposal cost, reflecting buyers’ stronger bargaining power and the reduced competitive pressure on resource demand.
(a) $s=2$, $c_{d}=10$
[…]
Mean transaction price trajectory for high scarcity ($s=2$) and low disposal cost ($c_{d}=10$).
(b) $s=0.5$, $c_{d}=20$
[…]
Mean transaction price trajectory for low scarcity ($s=0.5$) and high disposal cost ($c_{d}=20$).
Figure 2. Evolution of mean transaction prices under different scarcity and disposal cost conditions.
To better understand the efficiency and stability of the sellers’ adaptive strategies, we next examine the dynamics at the level of individual firms using counterfactual regret analysis zinkevich2007regret, which quantifies how much a seller could have improved its payoff by choosing alternative actions.
Let $\phi_{j}^{(t)}\in\Phi_{j}$ denote the action used by seller $s$ at timestep $t$, and $r_{j}^{(t)}(\phi_{j}^{(t)})$ denote the reward received at that step. The optimal reward at time $t$ is
$r_{j}^{\star(t)}=\max_{\phi\in\Phi_{j}}r_{j}^{(t)}(\phi),$ (7)
where $r_{j}^{(t)}(\phi)$ is the reward seller $s$ would have obtained at step $t$ if it had chosen action $\phi$ while all other sellers’ actions remained unchanged. The regret of seller $s$ at time $t$ is then
$\text{Regret}_{j}^{(t)}=r_{j}^{\star(t)}-r_{j}^{(t)}\left(\phi_{j}^{(t)}\right).$ (8)
A positive regret indicates that the seller could have achieved a higher payoff at that step by selecting a different action, whereas a regret of zero indicates that the chosen action was optimal given the choices of the other sellers at that moment.
[…]
Rolling median of per-step regret for high scarcity ($s=2$) and low disposal cost ($c_{d}=10$).
Figure 3. Rolling median ($window=50$) of per-step regret under different scarcity and disposal cost conditions.
Empirically, we observe that per-step total regret, the aggregate deviation from a joint optimum, declines over time and converges toward zero in both scenarios (Figure 3). Since per-step regret measures the immediate payoff loss relative to the best fixed action in hindsight at that step, a vanishing regret implies that sellers’ strategies are nearly optimal given the current actions of others. Consequently, the joint strategy profile of all sellers appears to converge toward an approximate Nash equilibrium, providing justification for the stability of the emergent market dynamics under our MARL learning scheme.
5.2. Price and Symbiosis Dynamics
In this section, we examine how firm density, resource scarcity, and disposal costs shape the dynamics of transaction prices and circularity. We focus on these output metrics because they directly quantify the level of local circularity achieved in the market and the average exchange price, making them relevant proxies for assessing the effectiveness of different policy levers. By systematically varying firm density, resource scarcity, and disposal costs, we analyze how structural and economic factors influence sellers’ adaptive strategies, shaping both the emergence of local symbiosis and overall market outcomes. Figures 4 and 5 illustrate the evolution of equilibrium prices and local symbiosis levels as a function of the disposal cost, the scarcity, and the density. Each point is the mean of 10 simulation runs, with shaded areas indicating one standard deviation to visualize variability across runs.
Figure 4. Evolution of the exchange price
Figure 5. Evolution of symbiosis index
On the price side (Figure 4), we observe that, as expected, higher disposal costs induce sellers to lower their prices in order to avoid the landfill penalty. This effect is strongest in low scarcity environments ($s=0.5$), where intense competition can even drive equilibrium prices below zero in dense territories. At intermediate scarcity ($s=1$), the downward trend remains but is attenuated, as sellers retain partial market power. Under high scarcity ($s=2$), prices remain largely insensitive to $c_{d}$: the excess demand ensures that buyers continue to pay close to the outside option $p_{m}$, leaving sellers with little incentive to adjust their strategies. Interestingly, at the demand equilibrium ($s=1$), we observe higher variance, reflecting the instability and potential oscillations in companies’ strategies as they adapt to competitive pressures.
Symbiosis levels (Figure 5) exhibit a complementary pattern: the share of demand met through local exchanges increases with $c_{d}$, as the disposal penalty strengthens the incentive to form local contracts. Here, density plays a critical role: in dense environments, local exchanges dominate rapidly and symbiosis saturates close to one; in sparse environments, by contrast, symbiosis remains low even under high $c_{d}$, reflecting structural limitations in firms’ connectivity. We observe that this time scarcity has only a marginal effect on the evolution of symbiosis: higher scarcity slightly enhances local exchanges, but the dominant factors remain disposal cost and firm density.
5.3. Sensitivity Analysis
The previous analysis showed that market outcomes are shaped by interacting mechanisms, but it did not allow us to fully disentangle their relative influence. To address this limitation and account for the influence of additional variables, we conducted a variance-based global sensitivity analysis saltelli2010variance. We evaluate the influence of the simulation parameters on the symbiosis index and converged mean price measured at the end of each simulation run, after 1000 timesteps. By analysing which parameters exert the greatest influence on the price and symbiosis index, we can identify the factors that policymakers should prioritize when designing interventions to foster circularity while keeping prices at a low level. We simulate for varying values of the five following variables:
  • Disposal cost ($c_{d}\in[0,200]$), representing the penalty for unsold byproducts;
  • Scarcity level ($s\in[0.25,2]$), controlling the relative abundancy of resources in the market;
  • Firm density ($\rho\in[10^{-5},10^{-1}]$), defining the number of firms per unit area;
  • Cluster spread ($cs\in[0,0.5]$), determining how tightly companies are spatially grouped within clusters;
  • Transportation cost per kilometer ($c_{t}\in[0,10]$), capturing the economic friction associated with distance.
To efficiently explore the parameter space, we employed surrogate models, which act as fast emulators of the full simulation while capturing the key input-output relationships saves2024smt. Specifically, we implemented a sparse Polynomial Chaos Expansion (PCE), enabling the analytical derivation of Sobol’ indices from its coefficients sudret2008global. The surrogate models demonstrated strong predictive performance, achieving an $R^{2}$ of 95% for the symbiosis and 92% for the price on a 20% test sample. Parameter sampling was performed using Latin hypercube sampling to train surrogate models, with $s$ sampled on a $\log_{2}$ scale and density{density} on a $\log_{10}$ scale. For each parameter set, two independent simulations were executed to account for stochastic variability in agent decisions, resulting in a total of 20,000 independent simulations.
Sobol’ sensitivity analysis decomposes the variance of the model output into orthogonal contributions attributable to individual input parameters and their interactions. We computed the full set of Sobol’ indices: first-order indices quantify the direct effect of each parameter on output variance, while higher-order indices capture interaction effects amongst parameters chastaing. Taken together, Sobol’ indices provide a rigorous and comprehensive assessment of the relative importance of independent inputs and their interactions, with their sum accounting for $100\%$ of the variance.
In Table 1, out of 31, we show only the 12 indices greater than $1\%$ of the total variance for either symbiosis or price indices. The Sobol’ analysis confirms the strong influence of firm density: higher density reduces the average distance between agents and intensifies competition, explaining its impact on both model outcomes. As expected, price is also strongly affected by scarcity, and both outputs are influenced by transport costs. Interestingly, higher-order interactions contribute substantially to the variance, emphasizing the importance of interplay effects between parameters.
Table 1. Comparison of Sobol’ sensitivity indices for the symbiosis and price outputs, including unaccounted variance from Sobol’ indices ¡ 1%.
Variables Price Sobol’ Symbiosis Sobol’
[…]
Even if higher density is highly influential for both our symbiosis and price objectives, in practice, density is constrained by the considered territory and is not directly actionable. Therefore, in the subsequent analysis, we focus on the 4 other input variables under different fixed density scenarios, since their interaction effects with density are all significant even at a third order. To better capture the non-linear effects, we replace the sparse PCE by a Multi-Layer Perceptron neural network surrogate model ($R^{2}$ of 97% for the symbiosis and 95% for the price), and use it to predict both the Partial Dependence Plots (PDP) and Individual Conditional Expectations (ICE) robani2025smt as illustrated in Figure 6. In these plots, each colored line corresponds to an ICE, showing how the outcome of a single simulation responds when varying one parameter while holding others fixed; the black line is the PDP, representing the average effect across all simulations, and the color of each ICE line encodes the underlying firm density.
First, for the price in Figure 6(a), the PDP and ICE plots are consistent with the Sobol’ analysis in Table 1. In the blue low-density cases, the exchange price remains close to the market price and is only weakly influenced by other variables. As density increases, however, the price tends to decrease. The most influential variable is Scarcity, which exhibits a threshold effect: at low levels, scarcity has little impact, but between 1.5 and 2.5 it strongly drives the price upward. Beyond 2.5, the effect persists but grows at a slower rate. We also observe that increasing the disposal cost provides a strong and linear incentive for price reduction; the higher the density, the stronger the effect. Finally, cluster spread and kilometer cost have more limited influence, suggesting that they are less effective levers for intervention on the price outcome.
Second, for circularity in Figure 6(b), we again observe a pronounced density effect: higher density leads to greater symbiosis, largely independent of other variables. At low density levels, cluster spread emerges as an important factor, but only when it takes on small values. As spread increases, its influence fades rapidly, suggesting that spatial compactness can partially compensate for low density but only within narrow limits. Reducing travel costs appears more consistently actionable: at low density, for example, a per-kilometer cost below $2$ is required to reach any desirable symbiosis value. Under the same low density conditions, increasing disposal costs also provide a clear incentive to exchange byproducts, while at high density its effect is only marginal. Finally, scarcity has a notable effect on low-to-mid density values only above the critical threshold of 2.5.
(a) PDP and ICE for the price index
(b) PDP and ICE for the symbiosis index
Figure 6. PDP and ICE of $c_{d},\ s,\ cd$ and $c_{t}$ colored by fixed densities for price and symbiosis.
In summary, firm density is a primary determinant of both symbiosis and price, though it is mostly exogenous and difficult to influence directly through policy. Amongst the controllable parameters, transport cost ($c_{t}$) and disposal penalty ($c_{d}$) emerge as effective levers. The sensitivity results indicate that reducing per-kilometer transport costs, through measures such as shared logistics or targeted subsidies, can substantially enhance exchange opportunities in low-density areas. Similarly, higher disposal penalties consistently incentivize firms to seek symbiotic exchanges instead of wasteful disposal.
6. Discussion
Our simulations highlight several insights into the dynamics of decentralized byproduct markets. First, the emergent pricing behavior shows that the RL process enables sellers to adapt efficiently to local conditions. The counterfactual regret analysis further demonstrates that the learned strategies converge toward a near Nash equilibrium, suggesting that decentralized adaptive learning can yield stable and rational outcomes without requiring a centralized market clearing mechanism. This finding underscores the potential of bottom-up coordination mechanisms in fostering circular exchanges.
The symbiosis index provides a tractable measure of local circularity and illustrates how territorial factors, scarcity, kilometer and disposal costs shape resource exchanges. Our results show that transportation costs and disposal penalties represent actionable levers, especially where density is not naturally high. Moreover, interactions among variables, including up to third-order effects, highlight that circularity outcomes result from the combined influence of spatial configuration and economic incentives rather than any single factor. These findings emphasize the importance of integrated, multivariate approaches when designing circular economy interventions.
Beyond methodological insights, the model offers a structured framework for evaluating policy instruments. The parameters of the simulation naturally correspond to concrete levers available to regulators. Increasing the disposal cost $c_{d}$ can be interpreted as a landfill tax, designed to encourage firms to recover or valorize byproducts locally instead of discarding them. Adjusting the external market price $p_{m}$ through tariffs or subsidies can favor local sourcing over reliance on virgin raw materials. Similarly, incorporating carbon related transportation costs km_costkm\_cost provides a means to internalize the environmental footprint of long distance exchanges. Finally, promoting spatial proximity through industrial zoning or cluster based planning can strengthen firm density, thereby enhancing the feasibility of symbiotic interactions.
While these policy-relevant insights highlight the practical applicability of our model, it is important to recognize several limitations that temper the generalizability of the results. The present model focuses on a single byproduct, which allows us to isolate core mechanisms of decentralized exchange but leaves aside interdependencies across multiple byproduct markets. Real-world industrial symbiosis often involves multi-product complementarities, cascading exchanges, and competition between alternative uses of resources. Likewise, our experiments relied on stylized virtual geographies rather than empirically calibrated territories. While this approach ensured control and replicability, it abstracts away from infrastructural constraints such as road networks or administrative boundaries that can strongly affect exchange feasibility. Another simplification concerns the temporal dimension: in our simulations, companies participate in auctions independently of their production cycles, whereas in practice production schedules and storage constraints strongly influence availability and willingness to trade. Furthermore, our agents trade without any memory of past interactions, while in reality firms build relationships over time, developing trust, reliability, and long term contracting practices. Nonetheless, by abstracting from these complexities, the present work provides an indispensable first step toward systematically modeling spatially constrained markets for industrial symbiosis, upon which richer and more realistic frameworks can be built.
These limitations also point to promising directions for future work. Extending the model to multi-product environments would allow capturing substitution effects and synergies between resource flows. Introducing temporal dynamics and agent memory would bring the model closer to real-world conditions, enabling the study of production rhythms, storage capacities, and relational contracts. Grounding the simulations in empirical case studies of specific territories would make it possible to validate predictions and directly inform policy design. Ultimately, developing a panel of decision parameters that systematically map simulation outcomes to actionable policies could provide decision makers with a robust tool to design interventions that align individual incentives with collective circular economy objectives.
7. Conclusion
We introduced a decentralized multi-agent model for local byproduct exchanges, capturing how sellers adapt prices through RL, based on spatial and economic constraints. Simulations exhibit that adaptive decentralized strategies can produce stable, efficient, and circular market outcomes, as evidenced by low counterfactual regret and high symbiosis indices. Our results highlight two complementary contributions.
First, from a policy design perspective, the simulator is the first step toward a decision support tool. By linking agent-level parameters to macro-level outcomes such as the symbiosis index, the model allows policymakers to identify which levers most strongly affect circularity. This provides a controlled environment where different regulatory scenarios can be tested before real-world implementation. The emergent insights from our simulations help policymakers prioritize interventions that effectively promote circular practices across heterogeneous firms and contexts.
Second, from a MAS perspective, the work provides a new methodological understanding of decentralized exchange mechanisms under spatial and economic constraints. We show how RL agents, interacting in a double auction market with heterogeneous resources and transport costs, can converge to stable trading patterns. Beyond the application domain, our contribution lies in revealing how market based coordination and learning dynamics shape emergent equilibria.
{acks}
This work is part of the MUTTEC project, funded by the France 2030 program – TIRIS (Interdisciplinarité et impact de la science pour la société), under contract number 23-AAP-TIRIS-01-062. This work is part of the MIMICO research project funded by the Agence Nationale de la Recherche (ANR) no ANR-24-CE23-0380.

### v2

@@
 Key words and phrases:
-Agent Based Simulation, Industrial Symbiosis, Circular Economy, Market Analysis
+Spatial Agent-Based Simulation, Industrial Symbiosis, Circular Economy, Distributed Market Analysis, Reinforcement Learning
+doi:
+EXII2056
 1. Introduction
@@
 Poor coordination often results in wasted resources and missed opportunities, highlighting the gap between individual strategies and collective circularity goals.
-Identifying the conditions under which decentralized firms can coordinate effectively is therefore a key research challenge.
+Identifying the conditions under which decentralized firms can coordinate effectively is, therefore, a key research challenge.
 Analytical solutions are often intractable due to the combinatorial complexity of interactions and the emergent nature of prices and traded quantities.
@@
 Agent-based models, in particular, are able to capture heterogeneity among actors and test how different market designs, decision rules, and learning mechanisms affect the emergence of exchanges under realistic economic constraints epstein1996growing.
-Their ability to explicitly represent geosituated agents makes them especially suitable for studying spatially dependent interactions, such as proximity based collaborations, transportation costs, and localized resource flows.
+Their ability to explicitly represent geosituated agents makes them especially suitable for studying spatially dependent interactions, such as proximity-based collaborations, transportation costs, and localized resource flows.
 Unlike theoretical game-theoretic models, which often rely on a small number of representative players to ensure tractability, agent-based models can accommodate many heterogeneous actors and interactions.
-However, their effectiveness is often constrained by the simplifications made in representing agent behaviors.
+i However, their effectiveness is often constrained by the simplifications made in representing agent behaviors.
 In many existing symbiosis models Yazan2020; Ghali2017; Raimbault2020; Gan2015; FernandezMena2020; Chahla2019, firms follow fixed behavioral rules or static pricing strategies, limiting their ability to capture the adaptive and uncertain nature of real-world exchanges.
@@
 A key feature of the model is that both transaction prices and traded quantities emerge endogenously from these interactions, allowing us to investigate how local decisions give rise to systemic patterns.
+While illustrated here through industrial symbiosis, this framework applies broadly to circular economy settings, such as heat recovery, wastewater reuse, or biomass valorization, where decentralized agents, ranging from firms to aggregated entities like neighborhoods, must coordinate to close resource loops under transport frictions.
 Our contributions are threefold.
 First, we develop a reinforcement learning-based market model for decentralized byproduct exchanges, based on a novel auction mechanism adapted to spatial and economic constraints of territorial circularity.
-Second, we validate the learned strategies via counterfactual regret analysis, showing convergence toward near-Nash equilibria.
-Finally, a simulation-based analysis of conditions under which adaptive pricing and agent interactions lead to stable, efficient, and circular outcomes, providing insights for policymakers.
+Second, we validate the learned strategies via counterfactual regret analysis, showing convergence toward near-Nash equilibrium.
+Finally, we propose a simulation-based analysis of the conditions under which adaptive pricing and agent interactions lead to stable circular outcomes, providing insights for policymakers.
 The paper first reviews related work (Section 2), then presents the decentralized market model (Section 3) and simulation setup (Section 4).
@@
 Yet, despite these advances, the explicit representation of economic incentives and market coordination mechanisms remains underexplored.
-Also, these works does not explicitly model markets where prices, trade volumes, and resource allocations emerge endogenously from decentralized interactions.
+Additionally, these works do not explicitly model markets where prices, trade volumes, and resource allocations emerge endogenously from decentralized interactions.
 2.2. Auctions and Market Mechanisms in Spatial Settings
@@
 Yet these contributions remain largely theoretical, typically involving few participants and limited attention to adaptive mechanisms that capture the evolving dynamics of real markets.
-In particular, the integration of spatial frictions with learning-based bidding strategies has received limited attention, despite its relevance for understanding decentralized coordination in resource-constrained environments.
 2.3. Learning Agents for Adaptive Market Participation
@@
 Reinforcement Learning (RL) overcomes these limitations by allowing agents to learn optimal policies through trial-and-error, balancing exploration and exploitation in complex, evolving environments.
-In the context of double auctions, RL agents are particularly well suited due to the complex interactions between buyers and sellers and the continuous feedback provided by order book dynamics.
+In the context of double auctions, RL agents are particularly well-suited due to the complex interactions between buyers and sellers and the continuous feedback provided by order book dynamics.
 Agents can be designed to adopt specific roles, such as deciding when to buy, sell, or set prices, enabling cooperative and strategic trading behaviors bjerkoey2019; maeda2020.
@@
 Yet, such constraints can be critical in markets where transportation or delivery costs influence trading decisions rodrigue2020geography.
-RL with spatial considerations have been primarily explored in logistics, where carriers and shippers act as bidders in dynamic auctions mediated by centralized brokers Wang2018; Qiao2019; zha2017surge; atasoy2020.
+RL with spatial considerations has been primarily explored in logistics, where carriers and shippers act as bidders in dynamic auctions mediated by centralized brokers Wang2018; Qiao2019; zha2017surge; atasoy2020.
 Even when shippers themselves are modeled as learning agents VanHeeswijk2019, exchanges remain intermediary mediated.
@@
   1. (1)
-    Industrial symbiosis models rarely provide in depth mechanism for endogenous market-based price formation.
+    Industrial symbiosis models rarely provide an in-depth mechanism for endogenous market-based price formation.
   2. (2)
@@
   3. (3)
-    Reinforcement learning in markets has not been applied to decentralized settings where spatial topology influence agents behavior.
-Our contribution is to bridge these strands by introducing a decentralized auction model for industrial symbiosis where (i) sellers adapt prices through reinforcement learning, (ii) spatial constraints are explicitly internalized, and (iii) both prices and traded quantities emerge from multilateral interactions.
-This framework connects industrial symbiosis modeling with mechanism design and multi-agent RL, allowing modelers to explore how local adaptive dynamics shape circularity outcomes.
+    Reinforcement learning in markets has not been applied to decentralized settings where spatial topology influences agents’ behavior.
+Our contribution is to bridge these strands by introducing a decentralized auction model for industrial symbiosis.
+While RL and double auctions have been studied in isolation, their integration into a spatial multilateral double-auction, where transport costs and location directly drive strategic behavior, constitutes a novel contribution.
+Specifically, our model ensures that (i) sellers adapt prices through reinforcement learning, (ii) spatial constraints are explicitly internalized, and (iii) both prices and traded quantities emerge from multilateral interactions.
 Our objective is to reproduce the systemic complexity of inter-firm interactions within industrial symbiosis as a subset of the broader circular economy, providing a foundational basis for modeling circular economic dynamics.
+Beyond the specific application, the contribution lies in the rigorous methodology and implementation of this multi-agent RL framework, offering a reusable and generalizable approach for studying spatial market interactions.
 3. Model
@@
 To concentrate on the essential dynamics of the model, we restrict our analysis to the exchange of a single byproduct.
-This simplification enables us to isolate and assess the effects of multilateral interactions and sellers strategies, without additional complexity introduced by multiple interdependent byproducts.
+This simplification enables us to isolate and assess the effects of multilateral interactions and sellers’ strategies, without additional complexity introduced by multiple interdependent byproducts.
 We model a decentralized market for one byproduct, composed of buyers $\mathcal{B}=\{B_{1},\dots,B_{N_{B}}\}$ and sellers $\mathcal{S}=\{S_{1},\dots,S_{N_{S}}\}$.
@@
 This mechanism ensures that buyers only engage in transactions that meet their cost expectations relative to the external market.
+We adopt here a ”Small Open Economy”GULATI2013288 assumption: we hypothesize that local industrial park internal exchanges are too small to impact the global commodity market price.
+Consequently, $p_{m}$ is treated as a fixed exogenous parameter.
 Sellers are endowed with quantities $q_{j}$ and propose personalized prices to buyers, defined as
 $p_{ij}=\phi_{j}\ p_{m}+c_{ij},$ (1)
-The term $c_{ij}$ is originally defined as $c_{ij}=d_{ij}\ c_{t}$, where $c_{t}$ represents the transportation cost per kilometer (expressed in percentage of the market price).
+The term $c_{ij}$ is originally defined as $c_{ij}=d_{ij}\ c_{t}$, where $c_{t}$ represents the transportation cost per kilometer (expressed as a percentage of the market price).
 This term can be further extended to capture the total logistics cost required for the byproduct provided by the seller to become usable by the buyer, including handling, treatment, or adaptation costs.
-Finally, $\phi_{j}$ is an adaptive pricing parameter, that evolves during the learning phase.
-In addition to market and transportation costs, sellers face a penalty $c_{d}$ for unsold quantities (also expressed in percentage of the market price), representing the cost of sending residual flows to landfill and reflecting both economic and environmental impacts.
+Finally, $\phi_{j}$ is an adaptive pricing parameter that evolves during the learning phase.
+In addition to market and transportation costs, sellers face a penalty $c_{d}$ for unsold quantities (also expressed as a percentage of the market price), representing the cost of sending residual flows to landfill and reflecting both economic and environmental impacts.
 If a seller $j$ ends a timestep with $q_{j}^{\text{unsold}}$ units remaining, its reward is reduced by $q_{j}^{\text{unsold}}\ c_{d}$, imposing a direct financial penalty that encourages sellers to adjust pricing strategies to minimize waste while still aiming to maximize profit.
@@
 The first term corresponds to the net revenue from executed contracts, while the second term penalizes unsold inventory through the landfill cost $c_{d}$.
-This spatially defined market closely resembles a spatial Bertrand oligopoly hotelling_stability_1929, where multiple sellers compete by setting prices and buyers prefer lower-cost options.
+This spatially defined market closely resembles a spatial Bertrand oligopoly hotelling_stability_1929, where multiple sellers compete by setting prices, and buyers prefer lower-cost options.
 Unlike the classical Bertrand model, where identical products drive prices toward marginal cost, the inclusion of transport costs differentiates sellers spatially, allowing local price variation and creating room for adaptive strategies to emerge.
@@
 We model the interactions between buyers and sellers using a decentralized multilateral auction described in Algorithm 1.
-Each seller submits a set of bids to each buyers, where each bid specifies a price and quantity for a given product $p$, computed on the basis of the logistic costs required to deliver the product to the buyer’s location and the external market price.
+Each seller submits a set of bids to each buyer, where each bid specifies a price and quantity for a given product $p$, computed on the basis of the logistic costs required to deliver the product to the buyer’s location and the external market price.
 Buyers evaluate the incoming bids and select the one that maximizes their utility, subject to their price sensitivity $\beta_{i}$ and demand $q_{i}$.
@@
 1:Buyers $\mathcal{B}$, Sellers $\mathcal{S}$, product $p$
-2:while contracts has been made at the previous timestep do
+2:while contracts have been made at the previous timestep do
 3:  Seller bids: each $S_{j}\in\mathcal{S}$ submits a bid $(q_{j},p_{ij})$ to each $B_{i}\in\mathcal{B}$, where $q_{j}$ is the maximum quantity available for $S_{j}$, and $p_{ij}=\phi_{j}\ p_{m}+d_{ij}\ c_{t}$
-4:  Buyer choice: each $B_{i}$ evaluates the received bids, selects the most attractive one within price tolerance $\beta_{i}p_{m}$, and sends acceptance proposal to the corresponding seller
+4:  Buyer choice: each $B_{i}$ evaluates the received bids, selects the most attractive one within price tolerance $\beta_{i}p_{m}$, and sends an acceptance proposal to the corresponding seller
 5:  Seller allocation: each $S_{j}$ sorts accepted bids by expected profit per sold unit, confirms contracts until inventory is exhausted, updates sold quantity and profit
@@
 Here, $\alpha\in[0,1]$ is the smoothing factor, which balances the influence of past and recent observations.
-By prioritizing recent data, EMA allows rapid adaptation to changing conditions while remaining computationally efficient, requiring only constant-time updates per action.
-Although it does not offer formal convergence guarantees, EMA provides a practical compromise between adaptivity and scalability, and allows to maintain high learning efficiency morales2024exponential.
+By prioritizing recent data, EMA allows for rapid adaptation to changing conditions while remaining computationally efficient, requiring only constant-time updates per action.
+Although it does not offer formal convergence guarantees, EMA provides a practical compromise between adaptivity and scalability and allows for maintaining high learning efficiency morales2024exponential.
 Sellers select their next action using Boltzmann (softmax) sampling sutton1998reinforcement with a declining temperature $\tau_{t}$:
@@
 where $\tau_{0}$ sets the initial exploration level, decay controls the annealing rate, and $\tau_{\min}$ prevents premature convergence.
-This ensures exploration dominates early in learning, while the policy gradually becomes greedier over time.
+This ensures exploration dominates early in learning, while the policy gradually becomes more conservative over time.
 The subsequent action $\phi_{j}^{(t+1)}$ is drawn at random from the set of actions according to probabilities $P_{k}$.
@@
 We finally introduce a symbiosis index (SI) to quantify the efficiency of local byproduct exchanges. It is defined as:
-$\text{SI}=\frac{q_{\text{bought}}}{\min({q_{\text{toSell}},q_{\text{needed}}})},$
-where $q_{\text{bought}}$ denotes the quantity actually purchased, $q_{\text{toSell}}$ the quantity offered by the seller, and $q_{\text{needed}}$ the quantity required by the buyer.
+$\text{SI}=\frac{q_{\text{bought}}}{\min({q_{\text{toSell}},q_{\text{needed}}})},$ (7)
+where $q_{\text{bought}}$ denotes the total quantity actually purchased, $q_{\text{toSell}}$ the total quantity offered by the sellers, and $q_{\text{needed}}$ the total quantity required by the buyers.
 The denominator normalizes the indicator so that it cannot exceed 1, even when the sold quantity is greater than the buyer’s need.
@@
 The businesses interact through the auction mechanism described earlier, with offers, acceptances, and executed contracts dynamically updating inventories and demands.
-The framework tracks a variety of systemic metrics over time, including:
-  • Local prices and their evolution;
-  • Proportion of demand satisfied locally (symbiosis index);
-  • Network connectivity and trading patterns;
-  • Distance travelled by the sold byproducts.
+The framework tracks a variety of systemic metrics over time, including the local prices and their evolution, the proportion of demand satisfied locally (symbiosis index) and the trading patterns.
 These indicators form the backbone of our analysis, as they connect local agent decisions to measurable systemic dynamics.
@@
 Visualization tools provide spatial and temporal representations of market dynamics, price formation, and circularity indicators.
-The code implementation is available on GitLab [^1] https://github.com/AAMAS746/aamas2026-746.
-4.2. Experimental Conditions.
-For similar reasons behind our decision to focus on a single product, we refrain from using a real-world territory and instead construct virtual environments.
+The code implementation is open source [^1] https://github.com/matthieu-mastio/MASTIO.
+4.2. Experimental Conditions
+For similar reasons behind our decision to focus on a single product, we decided to focalize on virtual environments.
 This approach provides precise control over key parameters such as firm density, territorial extent, and spatial clustering, thereby enabling a systematic exploration of their influence on circularity indicators and overall market performance.
 One of the key factors shaping the potential for cooperation is the spatial dispersion of firms.
-We therefore define the cluster spread parameter ($cs$) as the standard deviation of companies positions around cluster centers, normalized by the environment width: small values lead to compact, well separated clusters, while setting it to $1$ produces a configuration comparable to a fully uniform distribution across the environment, effectively removing any clustering effect.
-Figure 1 shows the difference between a low (1(a)) and high (1(b)) spread value.
+We therefore define the cluster spread parameter ($cs$) as the standard deviation of companies’ positions around cluster centers, normalized by the environment width: small values lead to compact, well-separated clusters, while setting it to $1$ produces a configuration comparable to a fully uniform distribution across the environment, effectively removing any clustering effect.
+Figure 1(b) shows the difference between a low (1(a)) and high (1(b)) spread value.
 In our experiments, businesses were allocated into four spatial clusters.
@@
 Figure 1. Comparison of spatial agent distributions for different cs values.
-The results we obtain came from simulations conducted with a population of 40 firms.
+[…]
+Comparison of spatial agent distributions for different cluster spread values: 2 cases are displayed, the first one with cs = 0.05, showing 4 dense clusters; the second one with cs = 0.5, displaying an almost uniform distribution of agents in space.
+The results we obtained came from simulations conducted with a population of 40 firms, representing the typical scale of a standard EIP.
 The market price was fixed at $100$, and the transportation cost was set to $0.1$ per kilometer.
+This transport cost is a normalized parameter relative to the market price, calibrated to ensure that spatial distance acts as a meaningful economic friction without completely blocking trade, a choice further supported by the sensitivity analysis in Section 5.3.
 Sellers adjusted their strategies through Boltzmann exploration, with a temperature decay parameter set to $0.996$ and the action space discretized into $30$ bins.
@@
 Mean transaction price trajectory for low scarcity ($s=0.5$) and high disposal cost ($c_{d}=20$).
-Figure 2. Evolution of mean transaction prices under different scarcity and disposal cost conditions.
+Figure 2. Evolution of mean transaction prices under different scarcity and disposal cost conditions, where $s$ denotes the scarcity ratio and $c_{d}$ the disposal cost.
 To better understand the efficiency and stability of the sellers’ adaptive strategies, we next examine the dynamics at the level of individual firms using counterfactual regret analysis zinkevich2007regret, which quantifies how much a seller could have improved its payoff by choosing alternative actions.
-Let $\phi_{j}^{(t)}\in\Phi_{j}$ denote the action used by seller $s$ at timestep $t$, and $r_{j}^{(t)}(\phi_{j}^{(t)})$ denote the reward received at that step. The optimal reward at time $t$ is
+Let $\phi_{j}^{(t)}\in\Phi_{j}$ denote the action used by seller $j$ at timestep $t$, and $r_{j}^{(t)}(\phi_{j}^{(t)})$ denote the reward received at that step. The optimal reward at time $t$ is
 $r_{j}^{\star(t)}=\max_{\phi\in\Phi_{j}}r_{j}^{(t)}(\phi),$ (8)
-where $r_{j}^{(t)}(\phi)$ is the reward seller $s$ would have obtained at step $t$ if it had chosen action $\phi$ while all other sellers’ actions remained unchanged.
-The regret of seller $s$ at time $t$ is then
+where $r_{j}^{(t)}(\phi)$ is the reward that the seller $j$ would have obtained at step $t$ if it had chosen action $\phi$ while all other sellers’ actions remained unchanged.
+The regret of seller $j$ at time $t$ is then
 $\text{Regret}_{j}^{(t)}=r_{j}^{\star(t)}-r_{j}^{(t)}\left(\phi_{j}^{(t)}\right).$ (9)
@@
 Figure 3. Rolling median ($window=50$) of per-step regret under different scarcity and disposal cost conditions.
-Empirically, we observe that per-step total regret, the aggregate deviation from a joint optimum, declines over time and converges toward zero in both scenarios (Figure 3).
+Empirically, we observe that the per-step total regret, the aggregate deviation from a joint optimum, declines over time and converges toward zero in both scenarios (Figure 3).
 Since per-step regret measures the immediate payoff loss relative to the best fixed action in hindsight at that step, a vanishing regret implies that sellers’ strategies are nearly optimal given the current actions of others.
@@
 By systematically varying firm density, resource scarcity, and disposal costs, we analyze how structural and economic factors influence sellers’ adaptive strategies, shaping both the emergence of local symbiosis and overall market outcomes.
-Figures 4 and 5 illustrate the evolution of equilibrium prices and local symbiosis levels as a function of the disposal cost, the scarcity, and the density.
+Figures 4 and 5 illustrate the evolution of equilibrium prices and local symbiosis levels as a function of disposal cost, scarcity, and density.
 Each point is the mean of 10 simulation runs, with shaded areas indicating one standard deviation to visualize variability across runs.
 Figure 4. Evolution of the exchange price
+[…]
+Evolution of the exchange price
 Figure 5. Evolution of symbiosis index
+[…]
+Evolution of symbiosis index
 On the price side (Figure 4), we observe that, as expected, higher disposal costs induce sellers to lower their prices in order to avoid the landfill penalty.
@@
 Symbiosis levels (Figure 5) exhibit a complementary pattern: the share of demand met through local exchanges increases with $c_{d}$, as the disposal penalty strengthens the incentive to form local contracts.
-Here, density plays a critical role: in dense environments, local exchanges dominate rapidly and symbiosis saturates close to one; in sparse environments, by contrast, symbiosis remains low even under high $c_{d}$, reflecting structural limitations in firms’ connectivity.
+Here, density plays a critical role: in dense environments, local exchanges dominate rapidly, and symbiosis saturates close to one; in sparse environments, by contrast, symbiosis remains low even under high $c_{d}$, reflecting structural limitations in firms’ connectivity.
 We observe that this time scarcity has only a marginal effect on the evolution of symbiosis: higher scarcity slightly enhances local exchanges, but the dominant factors remain disposal cost and firm density.
@@
 To address this limitation and account for the influence of additional variables, we conducted a variance-based global sensitivity analysis saltelli2010variance.
-We evaluate the influence of the simulation parameters on the symbiosis index and converged mean price measured at the end of each simulation run, after 1000 timesteps.
-By analysing which parameters exert the greatest influence on the price and symbiosis index, we can identify the factors that policymakers should prioritize when designing interventions to foster circularity while keeping prices at a low level.
-We simulate for varying values of the five following variables:
+We evaluate the influence of the simulation parameters on the symbiosis index and exchanged prices measured at the end of each simulation run after 1000 timesteps.
+By analyzing which parameters exert the greatest influence on the price and symbiosis index, we can identify the factors that policymakers should prioritize when designing interventions to foster circularity while keeping prices at a low level.
+We simulate varying values for the following five variables:
   • Disposal cost ($c_{d}\in[0,200]$), representing the penalty for unsold byproducts;
-  • Scarcity level ($s\in[0.25,2]$), controlling the relative abundancy of resources in the market;
+  • Scarcity level ($s\in[0.25,2]$), controlling the relative abundance of resources in the market;
   • Firm density ($\rho\in[10^{-5},10^{-1}]$), defining the number of firms per unit area;
@@
 Sobol’ sensitivity analysis decomposes the variance of the model output into orthogonal contributions attributable to individual input parameters and their interactions.
-We computed the full set of Sobol’ indices: first-order indices quantify the direct effect of each parameter on output variance, while higher-order indices capture interaction effects amongst parameters chastaing.
+We computed the full set of Sobol’ indices: first-order indices quantify the direct effect of each parameter on output variance, while higher-order indices capture interaction effects among parameters chastaing.
 Taken together, Sobol’ indices provide a rigorous and comprehensive assessment of the relative importance of independent inputs and their interactions, with their sum accounting for $100\%$ of the variance.
@@
 […]
-Even if higher density is highly influential for both our symbiosis and price objectives, in practice, density is constrained by the considered territory and is not directly actionable.
+Although firm density emerges as the most influential factor for both symbiosis and price, its overwhelming effect tends to dominate the variance and obscure the individual contributions of the other parameters.
+To better isolate and interpret the impact of the potentially actionable economic and spatial levers, we therefore fix density at different levels and focus the subsequent analysis on the remaining four input variables.
 Therefore, in the subsequent analysis, we focus on the 4 other input variables under different fixed density scenarios, since their interaction effects with density are all significant even at a third order.
-To better capture the non-linear effects, we replace the sparse PCE by a Multi-Layer Perceptron neural network surrogate model ($R^{2}$ of 97% for the symbiosis and 95% for the price), and use it to predict both the Partial Dependence Plots (PDP) and Individual Conditional Expectations (ICE) robani2025smt as illustrated in Figure 6.
+To better capture the non-linear effects, we replace the sparse PCE by a Multi-Layer Perceptron neural network surrogate model ($R^{2}$ of 97% for the symbiosis and 95% for the price), and use it to predict both the Partial Dependence Plots (PDP) and Individual Conditional Expectations (ICE) robani2025smt as illustrated in Figure 5.3.
 In these plots, each colored line corresponds to an ICE, showing how the outcome of a single simulation responds when varying one parameter while holding others fixed; the black line is the PDP, representing the average effect across all simulations, and the color of each ICE line encodes the underlying firm density.
@@
 Finally, scarcity has a notable effect on low-to-mid density values only above the critical threshold of 2.5.
-(a) PDP and ICE for the price index
-(b) PDP and ICE for the symbiosis index
+(a) PDP and ICE illustrating the factors influencing the price index
+(b) PDP and ICE illustrating the factors influencing the symbiosis index
 Figure 6. PDP and ICE of $c_{d},\ s,\ cd$ and $c_{t}$ colored by fixed densities for price and symbiosis.
-In summary, firm density is a primary determinant of both symbiosis and price, though it is mostly exogenous and difficult to influence directly through policy.
-Amongst the controllable parameters, transport cost ($c_{t}$) and disposal penalty ($c_{d}$) emerge as effective levers.
-The sensitivity results indicate that reducing per-kilometer transport costs, through measures such as shared logistics or targeted subsidies, can substantially enhance exchange opportunities in low-density areas.
-Similarly, higher disposal penalties consistently incentivize firms to seek symbiotic exchanges instead of wasteful disposal.
+[…]
+[…]
 6. Discussion
-Our simulations highlight several insights into the dynamics of decentralized byproduct markets.
-First, the emergent pricing behavior shows that the RL process enables sellers to adapt efficiently to local conditions.
-The counterfactual regret analysis further demonstrates that the learned strategies converge toward a near Nash equilibrium, suggesting that decentralized adaptive learning can yield stable and rational outcomes without requiring a centralized market clearing mechanism.
-This finding underscores the potential of bottom-up coordination mechanisms in fostering circular exchanges.
-The symbiosis index provides a tractable measure of local circularity and illustrates how territorial factors, scarcity, kilometer and disposal costs shape resource exchanges.
-Our results show that transportation costs and disposal penalties represent actionable levers, especially where density is not naturally high.
-Moreover, interactions among variables, including up to third-order effects, highlight that circularity outcomes result from the combined influence of spatial configuration and economic incentives rather than any single factor.
-These findings emphasize the importance of integrated, multivariate approaches when designing circular economy interventions.
-Beyond methodological insights, the model offers a structured framework for evaluating policy instruments.
-The parameters of the simulation naturally correspond to concrete levers available to regulators.
-Increasing the disposal cost $c_{d}$ can be interpreted as a landfill tax, designed to encourage firms to recover or valorize byproducts locally instead of discarding them.
-Adjusting the external market price $p_{m}$ through tariffs or subsidies can favor local sourcing over reliance on virgin raw materials.
-Similarly, incorporating carbon related transportation costs km_costkm\_cost provides a means to internalize the environmental footprint of long distance exchanges.
-Finally, promoting spatial proximity through industrial zoning or cluster based planning can strengthen firm density, thereby enhancing the feasibility of symbiotic interactions.
-While these policy-relevant insights highlight the practical applicability of our model, it is important to recognize several limitations that temper the generalizability of the results.
-The present model focuses on a single byproduct, which allows us to isolate core mechanisms of decentralized exchange but leaves aside interdependencies across multiple byproduct markets.
-Real-world industrial symbiosis often involves multi-product complementarities, cascading exchanges, and competition between alternative uses of resources.
-Likewise, our experiments relied on stylized virtual geographies rather than empirically calibrated territories.
-While this approach ensured control and replicability, it abstracts away from infrastructural constraints such as road networks or administrative boundaries that can strongly affect exchange feasibility.
-Another simplification concerns the temporal dimension: in our simulations, companies participate in auctions independently of their production cycles, whereas in practice production schedules and storage constraints strongly influence availability and willingness to trade.
-Furthermore, our agents trade without any memory of past interactions, while in reality firms build relationships over time, developing trust, reliability, and long term contracting practices.
-Nonetheless, by abstracting from these complexities, the present work provides an indispensable first step toward systematically modeling spatially constrained markets for industrial symbiosis, upon which richer and more realistic frameworks can be built.
-These limitations also point to promising directions for future work.
-Extending the model to multi-product environments would allow capturing substitution effects and synergies between resource flows.
-Introducing temporal dynamics and agent memory would bring the model closer to real-world conditions, enabling the study of production rhythms, storage capacities, and relational contracts.
-Grounding the simulations in empirical case studies of specific territories would make it possible to validate predictions and directly inform policy design.
-Ultimately, developing a panel of decision parameters that systematically map simulation outcomes to actionable policies could provide decision makers with a robust tool to design interventions that align individual incentives with collective circular economy objectives.
+Our results demonstrate that the proposed mechanism generates stable, economically grounded behaviors.
+Emergent pricing patterns reveal that sellers adapt efficiently to local conditions, while counterfactual regret analysis confirms convergence toward a near-Nash equilibrium rather than stochastic oscillation.
+Furthermore, global sensitivity analysis establishes clear causal links between model parameters and market outcomes.
+Significant high-order interactions suggest that circularity arises from the synergy of spatial configuration and economic incentives, highlighting the necessity for integrated, multivariate policy frameworks.
+Ultimately, these findings indicate that decentralized adaptive learning produces rational outcomes without centralized clearing, validating the mechanism’s robustness.
+Beyond methodology, the model offers a framework for evaluating policy instruments by mapping simulation parameters to regulatory levers.
+For instance, increasing disposal costs mimics landfill taxes, while the cluster spread parameter can be interpreted as a proxy for industrial zoning or policies that promote firm density.
+This mapping enables the simulator to function as a virtual laboratory, allowing stakeholders to perform “what-if” analyzes and compare interventions before real-world deployment.
+While these insights highlight the applicability of our model, we must recognize several limitations that temper the generalizability of the results.
+We focused on a single byproduct to isolate the core mechanisms of decentralized exchange, abstracting away product heterogeneity and multi-market interdependence.
+Likewise, to ensure control and replicability, our experiments relied on stylized virtual geographies rather than empirically calibrated territories.
+We also abstract from temporal dynamics (production cycles, storage) and agent memory (trust, long-term contracts).
+Nonetheless, this work provides an indispensable first step toward systematically modeling spatially constrained markets for industrial symbiosis, upon which richer and more realistic frameworks can be built.
+Future work will address these limitations by introducing multi-product environments, temporal dynamics, and agent memory to capture substitution effects and relational contracts.
+Grounding simulations in empirical case studies will further validate predictions and refine the policy decision tool.
+Finally, we also plan to include rigorous runtime benchmarking to formally showcase the simulator’s scalability.
 7. Conclusion
 We introduced a decentralized multi-agent model for local byproduct exchanges, capturing how sellers adapt prices through RL, based on spatial and economic constraints.
-Simulations exhibit that adaptive decentralized strategies can produce stable, efficient, and circular market outcomes, as evidenced by low counterfactual regret and high symbiosis indices.
+Simulations exhibit that adaptive decentralized strategies can produce stable circular outcomes.
 Our results highlight two complementary contributions.
-First, from a policy design perspective, the simulator is the first step toward a decision support tool.
+First, from a policy design perspective, the simulator is the initial step toward a decision support tool.
 By linking agent-level parameters to macro-level outcomes such as the symbiosis index, the model allows policymakers to identify which levers most strongly affect circularity.
@@
 We show how RL agents, interacting in a double auction market with heterogeneous resources and transport costs, can converge to stable trading patterns.
-Beyond the application domain, our contribution lies in revealing how market based coordination and learning dynamics shape emergent equilibria.
+Beyond the application domain, our contribution lies in revealing how market based coordination and learning dynamics shape emergent equilibrium.
 {acks}
-This work is part of the MUTTEC project, funded by the France 2030 program – TIRIS (Interdisciplinarité et impact de la science pour la société), under contract number 23-AAP-TIRIS-01-062.
-This work is part of the MIMICO research project funded by the Agence Nationale de la Recherche (ANR) no ANR-24-CE23-0380.
+This work was supported by the MUTTEC project (France 2030 – TIRIS, contract 23-AAP-TIRIS-01-062) and by the MIMICO project (ANR-24-CE23-0380).